### v29

# A short computer program that computes in the limit a non-computable function from $\mathbb{N}$ to $\mathbb{N}$

Abstract
For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $\gamma(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\gamma:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\gamma_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\gamma(n)$. The author is not aware about computer programs of other authors that compute in the limit non-computable functions from $\mathbb{N}$ to $\mathbb{N}$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
Key words and phrases: computable function, eventual domination, limit-computable function, non-computable function, single-fold Diophantine representation. 2020 Mathematics Subject Classification: 03D20.
1 Predicate $\mathcal{K}$} of the current mathematical knowledge
In sections 14, $\mathcal{K}$} denotes both the predicate satisfied by every currently known theorem and the set of all currently known theorems. The set $\mathcal{K}$} is time-dependent. Observations 15 explain the definition of $\mathcal{K}$}. This definition is intentionally simple, which restricts the set $\mathcal{K}$}. Observations 14 hold at any time.
Observation 1.
Every written down theorem from past or present belongs to $\mathcal{K}$}.
Observation 2.
If $\Delta\in{\mathcal{K}}$, then every particular case of $\Delta$ belongs to $\mathcal{K}$}. For example,
$\{\emptyset\cup\{n\}=\{n\}:~n\in\mathbb{N}\}\subseteq{\mathcal{K}}$
Observation 3.
Since the Replacement Axiom is a scheme of axioms, only finitely many axioms of $ZFC$ belong to $\mathcal{K}$}.
Observation 4.
$\mathcal{K}$} is not deductively closed.
Proof.
There exists $n\in\mathbb{N}$ such that $\neg{\mathcal{K}}(\underbrace{(2\cdot 2=4)\wedge\ldots\wedge(2\cdot 2=4)}_{{\rm conjunction~of~}n+3{\rm~statements}})$. ∎
Observation 5.
Currently,
$(\neg{\mathcal{K}}((((((((9!)!)!)!)!)!)!)!{\rm-th~digit~of~\pi~is~even}))\wedge$
$(\neg{\mathcal{K}}((((((((9!)!)!)!)!)!)!)!{\rm-th~digit~of~\pi~is~odd}))$
Hence, $\mathcal{K}$} is not deductively closed now.
2 A program that computes in the limit a function $\gamma:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$
For $n\in\mathbb{N}$, let
$E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$
Theorem 1.
([8, p. 118]). There exists a limit-computable function $\gamma:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$.
We present an alternative proof of Theorem 1. For $n\in\mathbb{N}$, $\gamma(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The function $\gamma:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$, see [10]. The term "dominated" in the title of [10] means "eventually dominated". Flowchart 1 shows a semi-algorithm which computes $\gamma(n)$ in the limit, see [10].
Flowchart 1
A semi-algorithm which computes $\gamma(n)$ in the limit
Corollary 1.
For every $k\in\mathbb{N}$, the function $\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$ is uncomputable.
Proposition 1.
If $k\in\mathbb{N}$, then the statement "the function $\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$ is uncomputable" belongs to $\mathcal{K}$}.
Proof.
It follows from Corollary 1 and Observations 1 and 2. ∎
Flowchart 2 shows a simpler semi-algorithm which computes $\gamma(n)$ in the limit.
Flowchart 2
A simpler semi-algorithm which computes $\gamma(n)$ in the limit
Lemma 1.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 2 does not exceed the number printed by Flowchart 1.
Proof.
For every $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$,
$E_{n}\supseteq\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Lemma 2.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 1 does not exceed the number printed by Flowchart 2.
Proof.
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $\mathcal{S}$}, then $(a_{0},\ldots,a_{n})$ solves the system of equations
$\widetilde{\mathcal{S}}:=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Theorem 2.
For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the same number.
Proof.
It follows from Lemmas 1 and 2. ∎
MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. The following program in MuPAD implements the semi-algorithm shown in Flowchart 2.
input("Input a non-negative integer n",n):
m:=0:
while TRUE do
X:=combinat::cartesianProduct([s \$s=0..m] \$t=0..n):
Y:=[max(op(X[u])) \$u=1..(m+1)^(n+1)]:
for p from 1 to (m+1)^(n+1) do
for q from 1 to (m+1)^(n+1) do
v:=1:
for k from 1 to n+1 do
if 1=X[p][k] and 1<>X[q][k] then v:=0 end_if:
for i from 1 to n+1 do
for j from i to n+1 do
if X[p][i]+X[p][j]=X[p][k] and X[q][i]+X[q][j]<>X[q][k] then v:=0 end_if:
if X[p][i]*X[p][j]=X[p][k] and X[q][i]*X[q][j]<>X[q][k] then v:=0 end_if:
end_for:
end_for:
end_for:
if max(op(X[q]))<max(op(X[p])) and v=1 then Y[p]:=0 end_if:
end_for:
end_for:
print(max(op(Y))):
m:=m+1:
end_while:
The author is not aware about computer programs of other authors that compute in the limit non-computable functions from $\mathbb{N}$ to $\mathbb{N}$.
For $n\in\mathbb{N}$, $\theta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq\{x_{i}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. From [10] and Lemma 3 in [9], it follows that the function $\theta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$. A bit shorter program in MuPAD computes $\theta$ in the limit.
3 A program that computes in the limit a function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation
The Davis-Putnam-Robinson-Matiyasevich theorem states that every listable set ${\mathcal{M}}\subseteq{\mathbb{N}}^{n}$ $(n\in\mathbb{N}\setminus\{0\})$ has a Diophantine representation, that is
$(a_{1},\ldots,a_{n})\in{\mathcal{M}}\Longleftrightarrow\exists x_{1},\ldots,x_{m}\in\mathbb{N}~~W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ (R)
for some polynomial $W$ with integer coefficients, see [5]. The representation (R) is said to be single-fold, if for any $a_{1},\ldots,a_{n}\in\mathbb{N}$ the equation $W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ has at most one solution $(x_{1},\ldots,x_{m})\in{\mathbb{N}}^{m}$. The representation (R) is said to be finite-fold, if for any $a_{1},\ldots,a_{n}\in\mathbb{N}$ the equation $W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ has only finitely many solutions $(x_{1},\ldots,x_{m})\in{\mathbb{N}}^{m}$.
Conjecture 1.
([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
Conjecture 2.
([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a finite-fold Diophantine representation.
Let $\Phi$ denote the following statement: the function $\mathbb{N}\ni n\to 2^{n}\in\mathbb{N}$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. For $n\in\mathbb{N}$, let
$g(n)=\left\{\begin{array}[]{cl}2^{n},&{\rm if~}\Phi{\rm~holds}\\ \gamma(n),&{\rm otherwise}\end{array}\right$.
The function $g:\mathbb{N}\to\mathbb{N}$ is computable if and only if $\Phi$ holds. Currently,
$(\neg{\mathcal{K}}(\Phi))\wedge(\neg{\mathcal{K}}(\neg\Phi))\wedge(\neg{\mathcal{K}}(g~is~computable))\wedge(\neg{\mathcal{K}}(g~is~uncomputable))$
Let $\Psi$ denote the following statement: the function $\mathbb{N}\ni n\to 2^{n}\in\mathbb{N}$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a finite-fold Diophantine representation. For $n\in\mathbb{N}$, let
$h(n)=\left\{\begin{array}[]{cl}2^{n},&{\rm if~}\Psi{\rm~holds}\\ \gamma(n),&{\rm otherwise}\end{array}\right$.
The function $h:\mathbb{N}\to\mathbb{N}$ is computable if and only if $\Psi$ holds. Currently,
$(\neg{\mathcal{K}}(\Psi))\wedge(\neg{\mathcal{K}}(\neg\Psi))\wedge(\neg{\mathcal{K}}(h~is~computable))\wedge(\neg{\mathcal{K}}(h~is~uncomputable))$
Lemma 3.
The function $g$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The function $h$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a finite-fold Diophantine representation.
Proof.
It follows from Theorem 1. ∎
For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The computability of $\beta$ is unknown.
Theorem 3.
The function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
This is proved in [10]. Flowchart 3 shows a semi-algorithm which computes $\beta(n)$ in the limit, see [10].
Flowchart 3
A semi-algorithm which computes $\beta(n)$ in the limit
Flowchart 4 shows a simpler semi-algorithm which computes $\beta(n)$ in the limit.
Flowchart 4
A simpler semi-algorithm which computes $\beta(n)$ in the limit
Lemma 4.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 4 does not exceed the number printed by Flowchart 3.
Proof.
For every $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$,
$E_{n}\supseteq\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
Lemma 5.
For every $n,m\in\mathbb{N}$, the number printed by Flowchart 3 does not exceed the number printed by Flowchart 4.
Proof.
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $\mathcal{S}$} in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system of equations
$\widehat{\mathcal{S}}:=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
$\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
$\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
By this and the inclusion $\widehat{\mathcal{S}}\supseteq{\mathcal{S}}$, $\widehat{\mathcal{S}}$ has exactly one solution in $\{0,\ldots,m\}^{n+1}$, namely $(a_{0},\ldots,a_{n})$. ∎
Theorem 4.
For every $n,m\in\mathbb{N}$, Flowcharts 3 and 4 print the same number.
Proof.
It follows from Lemmas 4 and 5. ∎
The following program in MuPAD implements the semi-algorithm shown in Flowchart 4.
input("Input a non-negative integer n",n):
m:=0:
while TRUE do
X:=combinat::cartesianProduct([s \$s=0..m] \$t=0..n):
Y:=[max(op(X[u])) \$u=1..(m+1)^(n+1)]:
for p from 1 to (m+1)^(n+1) do
for q from 1 to (m+1)^(n+1) do
v:=1:
for k from 1 to n+1 do
if 1=X[p][k] and 1<>X[q][k] then v:=0 end_if:
for i from 1 to n+1 do
for j from i to n+1 do
if X[p][i]+X[p][j]=X[p][k] and X[q][i]+X[q][j]<>X[q][k] then v:=0 end_if:
if X[p][i]*X[p][j]=X[p][k] and X[q][i]*X[q][j]<>X[q][k] then v:=0 end_if:
end_for:
end_for:
end_for:
if q<>p and v=1 then Y[p]:=0 end_if:
end_for:
end_for:
print(max(op(Y))):
m:=m+1:
end_while:
Statement 1.
There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
Proof.
Statement 1 follows constructively from Theorem 3 by taking $f=\beta$ because the following conjunction
$({\neg}{\mathcal{K}}(\beta{\rm~is~computable}))~\wedge~(\neg{\mathcal{K}}(\beta{\rm~is~uncomputable}))$
holds. Statement 1 follows non-constructively from Lemma 3 by taking $f=g$ because the following conjunction
$({\neg}{\mathcal{K}}(g{\rm~is~computable}))~\wedge~({\neg}{\mathcal{K}}(g{\rm~is~uncomputable}))$
holds. ∎
Since the function $\gamma$ in Theorem 1 is not computable, Statement 1 does not follow from Theorem 1.
Statement 2.
In Statement 1, we can require that there exists a computer program which takes as input a non-negative integer $n$ and prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
Proof.
Any computer program that implements the semi-algorithm shown in Flowchart 3 or 4 is correct. ∎
Proposition 2.
Statement 1 strengthens a mathematical theorem. Statement 1 refers to the current mathematical knowledge and may be false in the future. Statement 1 does not express what is currently unproved in mathematics.
Proof.
Statement 1 strengthens Statement 1 without the epistemic condition. The weakened Statement 1 is a theorem which follows from Theorem 1. Statement 1 claims that some mathematically defined function $f:\mathbb{N}\to\mathbb{N}$ satisfies
($f$ is computable in the limit) $\wedge$ (${\neg}{\mathcal{K}}(f$ is computable)) $\wedge$ $({\neg}{\mathcal{K}}(f$ is uncomputable)) $\wedge$
($f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation)
Conjecture 1 disproves Statement 1. ∎
Statement 3.
Statement 1 holds for finite-fold Diophantine representations.
Proof.
It follows from Lemma 3 by taking $f=h$ because the following conjunction
$({\neg}{\mathcal{K}}(h{\rm~is~computable}))~\wedge~({\neg}{\mathcal{K}}(h{\rm~is~uncomputable}))$
holds. ∎
Statement 3 strengthens Statement 1. For Statement 3, there is no known computer program that computes $f$ in the limit.
4 A currently true statement $\bf G$} of the form $\exists f:\mathbb{N}\to\mathbb{N}$ of unknown computability $\ldots$", where $f$ is conjecturally uncomputable, $\bf G$} remains unproven when $f$ is computable, and $\bf G$} holds forever when $f$ is uncomputable
Let $\bf G$} denote the following statement: $\exists f:\mathbb{N}\to\mathbb{N}$ of unknown computability such that $f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation and there is a computer program which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
Proposition 3.
The statement $\bf G$} has all properties from the title of this section.
Proposition 4.
The statement $\bf G$} may hold when ${\mathcal{K}}(\beta{\rm~is~computable})$.
Proof.
For $n\in\mathbb{N}$, $\xi(n)$ denotes the smallest $b\in\mathbb{N}\setminus\{0\}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{Q}}^{n+1}$, then $\mathcal{S}$} has a solution which consists of rationals whose numerators and denominators belong to $\{-b,\ldots,b\}$. The function $\xi:\mathbb{N}\to\mathbb{N}$ is computable in the limit. We skip the proof, which is similar to the proof that $\gamma$ is computable in the limit. Currently,
$(\neg{\mathcal{K}}(\xi{\rm~is~computable}))\wedge(\neg{\mathcal{K}}(\xi{\rm~is~uncomputable}))$
At any future time, the conjunction
$(\neg{\mathcal{K}}(\xi{\rm~is~computable}))\wedge(\neg{\mathcal{K}}(\xi{\rm~is~uncomputable}))\wedge{\mathcal{K}}(\beta{\rm~is~computable})$
may hold and implies the statement $\bf G$} with $f=\xi+\beta$. ∎
Summarizing, the predicate $\mathcal{K}$} can strengthen existential mathematical statements when $\mathcal{K}$} is inserted after $\exists$ and refers to a part of the statement.
5 Predicate $\mathcal{K}$} of the written down mathematical knowledge
In this section, $\mathcal{K}$} denotes both the predicate satisfied by every written down theorem and the finite set of all written down theorems. It changes what is taken as known in mathematics.
Proposition 5.
Since $\mathcal{K}$} is finite, there exists $k\in\mathbb{N}$ such that the computability of the function
$\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$
is unknown. For this $k$, Statements 1 and 3 hold when $f(n)=k+\gamma(n)$.
Proposition 5 contradicts Proposition 1 with the right definition of $\mathcal{K}$}.

### v30

@@
-A short computer program that computes in the limit a non-computable function from $\mathbb{N}$ to $\mathbb{N}$
+A short program in MuPAD that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$
 Abstract
-For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$.
-For $n\in\mathbb{N}$, $\gamma(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
-The author proved earlier that the function $\gamma:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$.
-We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\gamma_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\gamma(n)$.
-The author is not aware about computer programs of other authors that compute in the limit non-computable functions from $\mathbb{N}$ to $\mathbb{N}$.
+It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable.
+Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit.
+For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k}$, $x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$.
+For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
+The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
+We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
 For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
@@
 We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
+2020 Mathematics Subject Classification: 03D20.
 Key words and phrases: computable function, eventual domination, limit-computable function, non-computable function, single-fold Diophantine representation.
-2020 Mathematics Subject Classification: 03D20.
-1 Predicate $\mathcal{K}$} of the current mathematical knowledge
-In sections 14, $\mathcal{K}$} denotes both the predicate satisfied by every currently known theorem and the set of all currently known theorems.
-The set $\mathcal{K}$} is time-dependent.
-Observations 15 explain the definition of $\mathcal{K}$}.
-This definition is intentionally simple, which restricts the set $\mathcal{K}$}.
-Observations 14 hold at any time.
+1 Introduction
+Definition 1.
+(cf.
+[9, pp.
+233–235]).
+A computation in the limit of a function $f:\mathbb{N}\to\mathbb{N}$ is a semi-algorithm which takes as input a non-negative integer $n$ and for every $m\in\mathbb{N}$ prints a non-negative integer $\xi(n,m)$ such that $\lim\limits_{m\to\infty}\xi(n,m)=f(n)$.
+It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable, see Theorem 1.
+Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit.
+In particular, this observation applies to the proof of Theorem 1 in [8], see also Observation 1.
 Observation 1.
-Every written down theorem from past or present belongs to $\mathcal{K}$}.
-Observation 2.
-If $\Delta\in{\mathcal{K}}$, then every particular case of $\Delta$ belongs to $\mathcal{K}$}. For example,
-$\{\emptyset\cup\{n\}=\{n\}:~n\in\mathbb{N}\}\subseteq{\mathcal{K}}$
-Observation 3.
-Since the Replacement Axiom is a scheme of axioms, only finitely many axioms of $ZFC$ belong to $\mathcal{K}$}.
-Observation 4.
-$\mathcal{K}$} is not deductively closed.
-Proof.
-There exists $n\in\mathbb{N}$ such that $\neg{\mathcal{K}}(\underbrace{(2\cdot 2=4)\wedge\ldots\wedge(2\cdot 2=4)}_{{\rm conjunction~of~}n+3{\rm~statements}})$. ∎
-Observation 5.
-Currently,
-$(\neg{\mathcal{K}}((((((((9!)!)!)!)!)!)!)!{\rm-th~digit~of~\pi~is~even}))\wedge$
-$(\neg{\mathcal{K}}((((((((9!)!)!)!)!)!)!)!{\rm-th~digit~of~\pi~is~odd}))$
-Hence, $\mathcal{K}$} is not deductively closed now.
-2 A program that computes in the limit a function $\gamma:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$
+Let $\varphi$ be a computable bijection from $\mathbb{N}$ to the set of all Diophantine equations. For $n\in\mathbb{N}$, let
+$\theta(n)=\left\{\begin{array}[]{cl}1,&\text{if~}\varphi(n)\text{~is~solvable~in~non-negative~integers}\\ 0,&\text{otherwise}\end{array}\right$.
+The function $\theta:\mathbb{N}\to\mathbb{N}$ is computable in the limit.
+A negative solution to Hilbert’s 10th problem implies that the function $\theta$ is not computable.
+There is no known $\varphi$ for which there exists a short computer program that computes $\theta$ in the limit.
+MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b.
+In this article, we present a short program in MuPAD that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
+2 A limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$
 For $n\in\mathbb{N}$, let
@@
 Theorem 1.
-([8, p. 118]). There exists a limit-computable function $\gamma:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$.
+([8, p. 118]). There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
 We present an alternative proof of Theorem 1.
-For $n\in\mathbb{N}$, $\gamma(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
-The function $\gamma:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$, see [10].
+For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
+The function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$, see [11].
 The term "dominated" in the title of [11] means "eventually dominated".
-Flowchart 1 shows a semi-algorithm which computes $\gamma(n)$ in the limit, see [10].
+Flowchart 1 shows a semi-algorithm which computes $f(n)$ in the limit, see [11].
 Flowchart 1
-A semi-algorithm which computes $\gamma(n)$ in the limit
-Corollary 1.
-For every $k\in\mathbb{N}$, the function $\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$ is uncomputable.
-Proposition 1.
-If $k\in\mathbb{N}$, then the statement "the function $\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$ is uncomputable" belongs to $\mathcal{K}$}.
-Proof.
-It follows from Corollary 1 and Observations 1 and 2. ∎
-Flowchart 2 shows a simpler semi-algorithm which computes $\gamma(n)$ in the limit.
+A semi-algorithm which computes $f(n)$ in the limit
+3 A short program in MuPAD that computes $f$ in the limit
+Flowchart 2 shows a simpler semi-algorithm which computes $f(n)$ in the limit.
 Flowchart 2
-A simpler semi-algorithm which computes $\gamma(n)$ in the limit
+A simpler semi-algorithm which computes $f(n)$ in the limit
 Lemma 1.
@@
 Let $n,m\in\mathbb{N}$.
-For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $\mathcal{S}$}, then $(a_{0},\ldots,a_{n})$ solves the system of equations
-$\widetilde{\mathcal{S}}:=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
+For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $\mathcal{S}$}, then $(a_{0},\ldots,a_{n})$ solves the following system of equations:
+$\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
 $\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
@@
 It follows from Lemmas 1 and 2. ∎
-MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. The following program in MuPAD implements the semi-algorithm shown in Flowchart 2.
+Definition 2.
+An approximation of a tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ is a tuple $(y_{0},\ldots,y_{n})\in{\mathbb{N}}^{n+1}$ such that
+$(\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}+x_{j}=x_{k}\Rightarrow y_{i}+y_{j}=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$
+Observation 2.
+There exists a set ${\mathcal{A}}(n)\subseteq{\mathbb{N}}^{n+1}$ such that
+${\rm card}({\mathcal{A}}(n))\leqslant 2^{\textstyle{\rm card}(E_{n})}=2^{\textstyle n+1+2\cdot(n+1)^{3}}$
+and every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in ${\mathcal{A}}(n)$.
+Observation 3.
+f(n)f(n)equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
+Observation 4.
+For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in\{0,\ldots,m\}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
+The following program in MuPAD implements the semi-algorithm shown in Flowchart 2.
 input("Input a non-negative integer n",n):
@@
 end_while:
-The author is not aware about computer programs of other authors that compute in the limit non-computable functions from $\mathbb{N}$ to $\mathbb{N}$.
-For $n\in\mathbb{N}$, $\theta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq\{x_{i}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
-From [10] and Lemma 3 in [9], it follows that the function $\theta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $\alpha:\mathbb{N}\to\mathbb{N}$.
-A bit shorter program in MuPAD computes $\theta$ in the limit.
-3 A program that computes in the limit a function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation
+For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq\{x_{i}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
+From [11] and Lemma 3 in [10], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
+A bit shorter program in MuPAD computes $h$ in the limit.
+4 A limit-computable function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation
 The Davis-Putnam-Robinson-Matiyasevich theorem states that every listable set ${\mathcal{M}}\subseteq{\mathbb{N}}^{n}$ $(n\in\mathbb{N}\setminus\{0\})$ has a Diophantine representation, that is
@@
 The representation (R) is said to be single-fold, if for any $a_{1},\ldots,a_{n}\in\mathbb{N}$ the equation $W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ has at most one solution $(x_{1},\ldots,x_{m})\in{\mathbb{N}}^{m}$.
-The representation (R) is said to be finite-fold, if for any $a_{1},\ldots,a_{n}\in\mathbb{N}$ the equation $W(a_{1},\ldots,a_{n},x_{1},\ldots,x_{m})=0$ has only finitely many solutions $(x_{1},\ldots,x_{m})\in{\mathbb{N}}^{m}$.
-Conjecture 1.
+Hypothesis 1.
 ([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
-Conjecture 2.
-([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a finite-fold Diophantine representation.
-Let $\Phi$ denote the following statement: the function $\mathbb{N}\ni n\to 2^{n}\in\mathbb{N}$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
-For $n\in\mathbb{N}$, let
-$g(n)=\left\{\begin{array}[]{cl}2^{n},&{\rm if~}\Phi{\rm~holds}\\ \gamma(n),&{\rm otherwise}\end{array}\right$.
-The function $g:\mathbb{N}\to\mathbb{N}$ is computable if and only if $\Phi$ holds. Currently,
-$(\neg{\mathcal{K}}(\Phi))\wedge(\neg{\mathcal{K}}(\neg\Phi))\wedge(\neg{\mathcal{K}}(g~is~computable))\wedge(\neg{\mathcal{K}}(g~is~uncomputable))$
-Let $\Psi$ denote the following statement: the function $\mathbb{N}\ni n\to 2^{n}\in\mathbb{N}$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a finite-fold Diophantine representation.
-For $n\in\mathbb{N}$, let
-$h(n)=\left\{\begin{array}[]{cl}2^{n},&{\rm if~}\Psi{\rm~holds}\\ \gamma(n),&{\rm otherwise}\end{array}\right$.
-The function $h:\mathbb{N}\to\mathbb{N}$ is computable if and only if $\Psi$ holds. Currently,
-$(\neg{\mathcal{K}}(\Psi))\wedge(\neg{\mathcal{K}}(\neg\Psi))\wedge(\neg{\mathcal{K}}(h~is~computable))\wedge(\neg{\mathcal{K}}(h~is~uncomputable))$
-Lemma 3.
-The function $g$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
-The function $h$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a finite-fold Diophantine representation.
-Proof.
-It follows from Theorem 1. ∎
 For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
@@
 A semi-algorithm which computes $\beta(n)$ in the limit
+5 A short program in MuPAD that computes $\beta$ in the limit
 Flowchart 4 shows a simpler semi-algorithm which computes $\beta(n)$ in the limit.
@@
 Let $n,m\in\mathbb{N}$.
-For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $\mathcal{S}$} in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system of equations
-$\widehat{\mathcal{S}}:=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
+For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $\mathcal{S}$} in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system $\widehat{\mathcal{S}}$, where
+$\widehat{\mathcal{S}}=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
 $\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
@@
 Proof.
-It follows from Lemmas 4 and 5. ∎
+It follows from Lemmas 3 and 4. ∎
 The following program in MuPAD implements the semi-algorithm shown in Flowchart 4.
@@
 end_while:
-Statement 1.
-There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
-Proof.
-Statement 1 follows constructively from Theorem 3 by taking $f=\beta$ because the following conjunction
-$({\neg}{\mathcal{K}}(\beta{\rm~is~computable}))~\wedge~(\neg{\mathcal{K}}(\beta{\rm~is~uncomputable}))$
-holds. Statement 1 follows non-constructively from Lemma 3 by taking $f=g$ because the following conjunction
-$({\neg}{\mathcal{K}}(g{\rm~is~computable}))~\wedge~({\neg}{\mathcal{K}}(g{\rm~is~uncomputable}))$
-holds. ∎
-Since the function $\gamma$ in Theorem 1 is not computable, Statement 1 does not follow from Theorem 1.
-Statement 2.
-In Statement 1, we can require that there exists a computer program which takes as input a non-negative integer $n$ and prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
-Proof.
-Any computer program that implements the semi-algorithm shown in Flowchart 3 or 4 is correct. ∎
-Proposition 2.
-Statement 1 strengthens a mathematical theorem.
-Statement 1 refers to the current mathematical knowledge and may be false in the future.
-Statement 1 does not express what is currently unproved in mathematics.
-Proof.
-Statement 1 strengthens Statement 1 without the epistemic condition.
-The weakened Statement 1 is a theorem which follows from Theorem 1.
-Statement 1 claims that some mathematically defined function $f:\mathbb{N}\to\mathbb{N}$ satisfies
-($f$ is computable in the limit) $\wedge$ (${\neg}{\mathcal{K}}(f$ is computable)) $\wedge$ $({\neg}{\mathcal{K}}(f$ is uncomputable)) $\wedge$
-($f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation)
-Conjecture 1 disproves Statement 1. ∎
-Statement 3.
-Statement 1 holds for finite-fold Diophantine representations.
-Proof.
-It follows from Lemma 3 by taking $f=h$ because the following conjunction
-$({\neg}{\mathcal{K}}(h{\rm~is~computable}))~\wedge~({\neg}{\mathcal{K}}(h{\rm~is~uncomputable}))$
-holds. ∎
-Statement 3 strengthens Statement 1. For Statement 3, there is no known computer program that computes $f$ in the limit.
-4 A currently true statement $\bf G$} of the form $\exists f:\mathbb{N}\to\mathbb{N}$ of unknown computability $\ldots$", where $f$ is conjecturally uncomputable, $\bf G$} remains unproven when $f$ is computable, and $\bf G$} holds forever when $f$ is uncomputable
-Let $\bf G$} denote the following statement: $\exists f:\mathbb{N}\to\mathbb{N}$ of unknown computability such that $f$ eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation and there is a computer program which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
-Proposition 3.
-The statement $\bf G$} has all properties from the title of this section.
-Proposition 4.
-The statement $\bf G$} may hold when ${\mathcal{K}}(\beta{\rm~is~computable})$.
-Proof.
-For $n\in\mathbb{N}$, $\xi(n)$ denotes the smallest $b\in\mathbb{N}\setminus\{0\}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{Q}}^{n+1}$, then $\mathcal{S}$} has a solution which consists of rationals whose numerators and denominators belong to $\{-b,\ldots,b\}$.
-The function $\xi:\mathbb{N}\to\mathbb{N}$ is computable in the limit.
-We skip the proof, which is similar to the proof that $\gamma$ is computable in the limit.
-Currently,
-$(\neg{\mathcal{K}}(\xi{\rm~is~computable}))\wedge(\neg{\mathcal{K}}(\xi{\rm~is~uncomputable}))$
-At any future time, the conjunction
-$(\neg{\mathcal{K}}(\xi{\rm~is~computable}))\wedge(\neg{\mathcal{K}}(\xi{\rm~is~uncomputable}))\wedge{\mathcal{K}}(\beta{\rm~is~computable})$
-may hold and implies the statement $\bf G$} with $f=\xi+\beta$. ∎
-Summarizing, the predicate $\mathcal{K}$} can strengthen existential mathematical statements when $\mathcal{K}$} is inserted after $\exists$ and refers to a part of the statement.
-5 Predicate $\mathcal{K}$} of the written down mathematical knowledge
-In this section, $\mathcal{K}$} denotes both the predicate satisfied by every written down theorem and the finite set of all written down theorems. It changes what is taken as known in mathematics.
-Proposition 5.
-Since $\mathcal{K}$} is finite, there exists $k\in\mathbb{N}$ such that the computability of the function
-$\mathbb{N}\ni n\to k+\gamma(n)\in\mathbb{N}$
-is unknown. For this $k$, Statements 1 and 3 hold when $f(n)=k+\gamma(n)$.
-Proposition 5 contradicts Proposition 1 with the right definition of $\mathcal{K}$}.

### v31

@@
-A short program in MuPAD that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$
+A short program that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$
 Abstract
@@
 Key words and phrases: computable function, eventual domination, limit-computable function, non-computable function, single-fold Diophantine representation.
-1 Introduction
+1 The Collatz problem leads to a short program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
 Definition 1.
 (cf.
-[9, pp.
+[10, pp.
 233–235]).
 A computation in the limit of a function $f:\mathbb{N}\to\mathbb{N}$ is a semi-algorithm which takes as input a non-negative integer $n$ and for every $m\in\mathbb{N}$ prints a non-negative integer $\xi(n,m)$ such that $\lim\limits_{m\to\infty}\xi(n,m)=f(n)$.
+By Definition 1, a function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit when there exists an infinite computation which takes as input a non-negative integer $n$ and prints a non-negative integer on each iteration and prints $f(n)$ on each sufficiently high iteration.
 It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable, see Theorem 1.
 Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit.
-In particular, this observation applies to the proof of Theorem 1 in [8], see also Observation 1.
-Observation 1.
-Let $\varphi$ be a computable bijection from $\mathbb{N}$ to the set of all Diophantine equations. For $n\in\mathbb{N}$, let
-$\theta(n)=\left\{\begin{array}[]{cl}1,&\text{if~}\varphi(n)\text{~is~solvable~in~non-negative~integers}\\ 0,&\text{otherwise}\end{array}\right$.
-The function $\theta:\mathbb{N}\to\mathbb{N}$ is computable in the limit.
-A negative solution to Hilbert’s 10th problem implies that the function $\theta$ is not computable.
-There is no known $\varphi$ for which there exists a short computer program that computes $\theta$ in the limit.
-MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b.
-In this article, we present a short program in MuPAD that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
+So far, short computer programs can only compute in the limit functions from $\mathbb{N}$ to $\mathbb{N}$ whose computability is proven or unknown.
+Lemma 1.
+For every $n\in\mathbb{N}$,
+$\frac{\textstyle\text{sign}(n-1)\cdot(2n+(1-(-1)^{n})\cdot(5n+2))}{\textstyle 4}=\left\{\begin{array}[]{rl}0{\rm,}&{\rm if~}n=1\\ \frac{\textstyle n}{\textstyle 2}{\rm,}&{\rm if~}n{\rm~is~even}\\ 3n+1{\rm,}&{\rm if~}n{\rm~is~odd~and~}n\neq 1\end{array}\right$.
+MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. By Lemma 1, the following program in MuPAD computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$.
+input("Input a non-negative integer n",n):
+while TRUE do
+print(sign(n)):
+n:=sign(n-1)*(2*n+(1-(-1)^n)*(5*n+2))/4:
+end_while:
+The computability of $\gamma$ is unknown, see [1, p. 79]. The Collatz conjecture implies that $\gamma(n)=0$ for every $n\in\mathbb{N}$.
 2 A limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$
@@
 Theorem 1.
-([8, p. 118]). There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
+([9, p. 118]). There exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
 We present an alternative proof of Theorem 1.
@@
 Proof.
-It follows from Lemmas 1 and 2. ∎
+It follows from Lemmas 2 and 3. ∎
 Definition 2.
@@
 $(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$
-Observation 2.
+Observation 1.
 There exists a set ${\mathcal{A}}(n)\subseteq{\mathbb{N}}^{n+1}$ such that
@@
 and every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in ${\mathcal{A}}(n)$.
+Observation 2.
+f(n)f(n)equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
 Observation 3.
-f(n)f(n)equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
-Observation 4.
-For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in\{0,\ldots,m\}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
+For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the smallest $b\in\{0,\ldots,m\}$ such that every tuple $(x_{0},\ldots,x_{n})\in\{0,\ldots,m\}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
 The following program in MuPAD implements the semi-algorithm shown in Flowchart 2.
@@
 Hypothesis 1.
-([1], [2], [3], [4, pp. 341–342], [6, p. 42], [7, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
+([2], [3], [4], [5, pp. 341–342], [7, p. 42], [8, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
 For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
@@
 Proof.
-It follows from Lemmas 3 and 4. ∎
+It follows from Lemmas 4 and 5. ∎
 The following program in MuPAD implements the semi-algorithm shown in Flowchart 4.

### v32

@@
-A short program that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$
+Two undecidable decision problems on an ordered pair of non-negative integers
 Abstract
-It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable.
-Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit.
-For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k}$, $x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$.
+For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$.
 For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
@@
 We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
+Since $f$ is not computable, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))\wedge$ $(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}+x_{j}=x_{k}\Rightarrow y_{i}+y_{j}=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow$ $y_{i}\cdot y_{j}=y_{k}))$.
+Similarly, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall~(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists~(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$.
 For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
@@
 We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
-2020 Mathematics Subject Classification: 03D20.
-Key words and phrases: computable function, eventual domination, limit-computable function, non-computable function, single-fold Diophantine representation.
-1 The Collatz problem leads to a short program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
+2020 Mathematics Subject Classification: 03D20, 11U05.
+Key words and phrases: computable function, eventual domination, limit-computable function, single-fold Diophantine representation, undecidable decision problem.
+1 The Collatz problem leads to a short computer program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
 Definition 1.
@@
 A semi-algorithm which computes $f(n)$ in the limit
-3 A short program in MuPAD that computes $f$ in the limit
+3 The first undecidable decision problem on an ordered pair of non-negative integers
 Flowchart 2 shows a simpler semi-algorithm which computes $f(n)$ in the limit.
@@
 Observation 1.
-There exists a set ${\mathcal{A}}(n)\subseteq{\mathbb{N}}^{n+1}$ such that
+For every $n\in\mathbb{N}$, there exists a set ${\mathcal{A}}(n)\subseteq{\mathbb{N}}^{n+1}$ such that
 ${\rm card}({\mathcal{A}}(n))\leqslant 2^{\textstyle{\rm card}(E_{n})}=2^{\textstyle n+1+2\cdot(n+1)^{3}}$
@@
 Observation 2.
-f(n)f(n)equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
+For every $n\in\mathbb{N}$, $f(n)$ equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
 Observation 3.
 For every $n,m\in\mathbb{N}$, Flowcharts 1 and 2 print the smallest $b\in\{0,\ldots,m\}$ such that every tuple $(x_{0},\ldots,x_{n})\in\{0,\ldots,m\}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
+Theorem 3.
+No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence: every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,m\}^{n+1}$.
+Proof.
+Since the function $f$ is not computable, it follows from Observation 2. ∎
+4 A short program in MuPAD that computes $f$ in the limit
 The following program in MuPAD implements the semi-algorithm shown in Flowchart 2.
@@
 end_while:
-For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq\{x_{i}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
+5 The second undecidable decision problem on an ordered pair of non-negative integers
+For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $\mathcal{S}$$\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
 From [12] and Lemma 3 in [11], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
 A bit shorter program in MuPAD computes $h$ in the limit.
+Theorem 4.
+No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence:
+$\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}$
+$(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$
+Proof.
+It holds because the function $h$ is not computable. ∎
 6 A limit-computable function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation

### v33

@@
-Two undecidable decision problems on an ordered pair of non-negative integers
+An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$
 Abstract
@@
 We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
-Since $f$ is not computable, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))\wedge$ $(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}+x_{j}=x_{k}\Rightarrow y_{i}+y_{j}=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow$ $y_{i}\cdot y_{j}=y_{k}))$.
-Similarly, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall~(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists~(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$.
+We prove that no algorithm takes as input a non-negative integer $n$ and decides whether or not $\exists p,q\in\mathbb{N}$ $((n=2^{p}\cdot 3^{q})$$\wedge$ $(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}~(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$.
 For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
@@
 2020 Mathematics Subject Classification: 03D20, 11U05.
-Key words and phrases: computable function, eventual domination, limit-computable function, single-fold Diophantine representation, undecidable decision problem.
+Key words and phrases: arithmetic of $\mathbb{N}$, computable function, eventual domination, limit-computable function, single-fold Diophantine representation, undecidable decision problem.
 1 The Collatz problem leads to a short computer program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
@@
 A semi-algorithm which computes $f(n)$ in the limit
-3 The first undecidable decision problem on an ordered pair of non-negative integers
+3 An undecidable decision problem about an ordered pair $(n,m)$ of non-negative integers that concerns the arithmetic of $\mathbb{N}$
 Flowchart 2 shows a simpler semi-algorithm which computes $f(n)$ in the limit.
@@
 end_while:
-5 The second undecidable decision problem on an ordered pair of non-negative integers
+5 An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$
 For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $\mathcal{S}$$\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
@@
 It holds because the function $h$ is not computable. ∎
+Corollary 1.
+No algorithm takes as input a non-negative integer $n$ and returns the logical value of the following sentence:
+$\exists p,q\in\mathbb{N}~((n=2^{p}\cdot 3^{q})\wedge$
+$(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$
+$(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$
 6 A limit-computable function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation

### v34

@@
-An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$
+Two undecidable decision problems about a non-negative integer $n$ which have a short description in terms of arithmetic
 Abstract
+We prove that the following decision problems (1) and (2) are algorithmically undecidable when $n\in\mathbb{N}$.
+(1) $\exists(y_{0},\ldots,y_{n})\in{\mathbb{N}}^{n+1}$ $\forall i,j,k\in\{0,\ldots,n\}$ $(((2^{\textstyle 2^{\textstyle 2^{\textstyle j}\cdot 3^{\textstyle k}}}+1$ divides $n)\Rightarrow$ $(y_{j}+1=y_{k}))$ $\wedge$ $((2^{\textstyle 2^{\textstyle 2^{\textstyle i}\cdot 3^{\textstyle j}\cdot 5^{\textstyle k+1}}}+1$ divides $n)$ $\Rightarrow$ $(y_{i}\cdot y_{j}=y_{k})))$.
+(2) $\exists p,q\in\mathbb{N}$ $((n=2^{p}\cdot 3^{q})$ $\wedge$ $\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}$ $\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$ $\forall i,j,k\in\{0,\ldots,p\}$ $(((x_{j}+1=x_{k})\Rightarrow(y_{j}+1=y_{k}))\wedge((x_{i}\cdot x_{j}=x_{k})\Rightarrow(y_{i}\cdot y_{j}=y_{k}))))$.
 For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$.
-For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
+For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $S$ has a solution in $\{0,\ldots,b\}^{n+1}$.
 The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
 We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$.
-We prove that no algorithm takes as input a non-negative integer $n$ and decides whether or not $\exists p,q\in\mathbb{N}$ $((n=2^{p}\cdot 3^{q})$$\wedge$ $(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}~(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$.
-For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
+For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
 The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation.
@@
 (cf.
-[10, pp.
+[11, pp.
 233–235]).
@@
 We present an alternative proof of Theorem 1.
-For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
+For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $S$ has a solution in $\{0,\ldots,b\}^{n+1}$.
 The function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$, see [13].
@@
 A semi-algorithm which computes $f(n)$ in the limit
-3 An undecidable decision problem about an ordered pair $(n,m)$ of non-negative integers that concerns the arithmetic of $\mathbb{N}$
+3 An undecidable decision problem about an ordered pair $(n,m)$ of non-negative integers which has a short description in terms of arithmetic
 Flowchart 2 shows a simpler semi-algorithm which computes $f(n)$ in the limit.
@@
 Let $n,m\in\mathbb{N}$.
-For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $\mathcal{S}$}, then $(a_{0},\ldots,a_{n})$ solves the following system of equations:
+For every system of equations $S\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $S$, then $(a_{0},\ldots,a_{n})$ solves the following system of equations:
 $\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
@@
 Theorem 3.
-No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence: every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,m\}^{n+1}$.
+No algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not
+$\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}$
+$((\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}+x_{j}=x_{k}\Rightarrow y_{i}+y_{j}=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k})))$
 Proof.
@@
 end_while:
-5 An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$
-For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $\mathcal{S}$$\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$.
-From [12] and Lemma 3 in [11], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
+5 Two undecidable decision problems about a non-negative integer $n$ which have a short description in terms of arithmetic
+Lemma 4.
+([12, p. 110]). For non-negative integers, the equation $x+y=z$ is equivalent to a system which consists of equations of the forms $\beta+1=\gamma$ and $\alpha\cdot\beta=\gamma$.
+For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $S$ has a solution in $\{0,\ldots,b\}^{n+1}$.
+From Lemma 4 and [13], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$.
 A bit shorter program in MuPAD computes $h$ in the limit.
 Theorem 4.
-No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence:
+No algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not
 $\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}$
-$(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
-$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$
+$((\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k})))$
 Proof.
@@
 Corollary 1.
-No algorithm takes as input a non-negative integer $n$ and returns the logical value of the following sentence:
+No algorithm takes as input a non-negative integer $n$ and decides whether or not
 $\exists p,q\in\mathbb{N}~((n=2^{p}\cdot 3^{q})\wedge$
-$(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$
-$(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
+$\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$
+$((\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
 $(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$
+Lemma 5.
+([10, p. 7]). For every $\in\mathbb{N}, xyx\neq y$ implies that the numbers $2^{\textstyle 2^{\textstyle x}}+1$ and $2^{\textstyle 2^{\textstyle y}}+1$ are relatively prime.
+Lemma 6.
+([12, p.
+110]).
+There exists a constructive algorithm that takes as input a Diophantine equation $D(x_{0},\ldots,x_{l})=0$ and returns a system $S$ of equations of the forms $y_{j}+1=y_{k}$ and $y_{i}\cdot y_{j}=y_{k}$ which is solvable in non-negative integers if and only if the equation $D(x_{0},\ldots,x_{l})=0$ is solvable in non-negative integers.
+Theorem 5.
+No algorithm takes as input a non-negative integer $n$ and decides whether or not
+$\exists(y_{0},\ldots,y_{n})\in{\mathbb{N}}^{n+1}~\forall i,j,k\in\{0,\ldots,n\}$ (P)
+$((2^{\textstyle 2^{\textstyle 2^{\textstyle j}\cdot 3^{\textstyle k}}}+1{\rm~divides~}n)\Rightarrow(y_{j}+1=y_{k}))\wedge((2^{\textstyle 2^{\textstyle 2^{\textstyle i}\cdot 3^{\textstyle j}\cdot 5^{\textstyle k+1}}}+1{\rm~divides~}n)\Rightarrow(y_{i}\cdot y_{j}=y_{k}))$
+Proof.
+If $n>0$, then we can compute a unique $(p,q)\in{\mathbb{N}}^{2}$ such that $n=2^{p}\cdot(2q+1)$.
+The decision problem (P) is algorithmically undecidable because we can obtain undecidability when $n>0$ and for every $i,j,k\in\{0,\ldots,n\}$
+$(\max(j,k)>p)\Rightarrow(2^{\textstyle 2^{\textstyle 2^{\textstyle j}\cdot 3^{\textstyle k}}}+1{\rm~does~not~divide~}n)$
+and
+$(\max(i,j,k)>p)\Rightarrow(2^{\textstyle 2^{\textstyle 2^{\textstyle i}\cdot 3^{\textstyle j}\cdot 5^{\textstyle k+1}}}+1{\rm~does~not~divide~}n)$
+In this case, by Lemma 5, for every system of equations
+$S\subseteq\{y_{j}+1=y_{k},~y_{i}\cdot y_{j}=y_{k}:~i,j,k\in\{0,\ldots,p\}\}$
+the problem of solvability of $S$ in non-negative integers $y_{0},\ldots,y_{p}$ is equivalent to the problem (P) for some $n=2^{p}\cdot(2q+1)$, where $n$ can be computed.
+Next, we apply Lemma 6 and a negative solution to Hilbert’s 10th problem.
+∎
 6 A limit-computable function $\beta:\mathbb{N}\to\mathbb{N}$ of unknown computability which eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation
@@
 ([2], [3], [4], [5, pp. 341–342], [7, p. 42], [8, p. 745]). Every listable set ${\mathcal{X}}\subseteq{\mathbb{N}}^{k}$ $(k\in\mathbb{N}\setminus\{0\})$ has a single-fold Diophantine representation.
-For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
+For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$.
 The computability of $\beta$ is unknown.
@@
 Let $n,m\in\mathbb{N}$.
-For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $\mathcal{S}$} in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system $\widehat{\mathcal{S}}$, where
-$\widehat{\mathcal{S}}=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
+For every system of equations $S\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $S$ in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system $\widehat{S}$, where
+$\widehat{S}=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
 $\{x_{i}+x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}+a_{j}=a_{k})\}\cup$
 $\{x_{i}\cdot x_{j}=x_{k}:~(i,j,k\in\{0,\ldots,n\})\wedge(a_{i}\cdot a_{j}=a_{k})\}$
-By this and the inclusion $\widehat{\mathcal{S}}\supseteq{\mathcal{S}}$, $\widehat{\mathcal{S}}$ has exactly one solution in $\{0,\ldots,m\}^{n+1}$, namely $(a_{0},\ldots,a_{n})$. ∎
+By this and the inclusion $\widehat{S}\supseteq S$, $\widehat{S}$ has exactly one solution in $\{0,\ldots,m\}^{n+1}$, namely $(a_{0},\ldots,a_{n})$. ∎
 Theorem 7.
@@
 Proof.
-It follows from Lemmas 4 and 5. ∎
+It follows from Lemmas 7 and 8. ∎
 The following program in MuPAD implements the semi-algorithm shown in Flowchart 4.

### v35

@@
 For every $n\in\mathbb{N}$,
-$\frac{\textstyle\text{sign}(n-1)\cdot(2n+(1-(-1)^{n})\cdot(5n+2))}{\textstyle 4}=\left\{\begin{array}[]{rl}0{\rm,}&{\rm if~}n=1\\ \frac{\textstyle n}{\textstyle 2}{\rm,}&{\rm if~}n{\rm~is~even}\\ 3n+1{\rm,}&{\rm if~}n{\rm~is~odd~and~}n\neq 1\end{array}\right$.
+$\frac{\textstyle\text{sgn}(n-1)\cdot(2n+(1-(-1)^{n})\cdot(5n+2))}{\textstyle 4}=\left\{\begin{array}[]{rl}0{\rm,}&{\rm if~}n=1\\ \frac{\textstyle n}{\textstyle 2}{\rm,}&{\rm if~}n{\rm~is~even}\\ 3n+1{\rm,}&{\rm if~}n{\rm~is~odd~and~}n\neq 1\end{array}\right$.
 MuPAD is a part of the Symbolic Math Toolbox in MATLAB R2019b. By Lemma 1, the following program in MuPAD computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$.
@@
 $(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$
+Let $\cdot$] denote the integer part function.
+Lemma 5.
+The function
+$\mathbb{N}\ni n\stackrel{{\scriptstyle{\textstyle\theta}}}{{{\textstyle\longrightarrow}}}(n-[\sqrt{n}]^{2},~|n-[\sqrt{n}]^{2}-[\sqrt{n}]|)\in{\mathbb{N}}^{2}$
+is surjective.
+Proof.
+It holds because $\theta(\{i^{2}+j:~(i,j\in\mathbb{N})\wedge(j\leqslant i)\})={\mathbb{N}}^{2}$. ∎
+Theorem 5.
+No algorithm takes as input a non-negative integer $n$ and decides whether or not
+$\forall(x_{0},\ldots,x_{n-[\sqrt{n}]^{2}})\in{\mathbb{N}}^{n-[\sqrt{n}]^{2}+1}~\exists(y_{0},\ldots,y_{n-[\sqrt{n}]^{2}})\in\{0,\ldots,|n-[\sqrt{n}]^{2}-[\sqrt{n}]|\}^{n-[\sqrt{n}]^{2}+1}$
+$((\forall j,k\in\{0,\ldots,n-[\sqrt{n}]^{2}\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
+$(\forall i,j,k\in\{0,\ldots,n-[\sqrt{n}]^{2}\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k})))$
+Proof.
+It follows from Theorem 4 and Lemma 5. ∎
 Lemma 6.
@@
 Proof.
-It follows from Lemmas 7 and 8. ∎
+It follows from Lemmas 8 and 9. ∎
 The following program in MuPAD implements the semi-algorithm shown in Flowchart 4.